Find the pH and percent ionization of each HF solution. (Ka for HF is 6.8×10^−4.)

1) Find the pH of a 0.250 M HF solution.
2) Find the percent dissociation of a 0.250 M HF solution.
3) Find the pH of a 0.130 M HF solution.
4) Find the percent dissociation of a 0.130 M HF solution.
5) Find the pH of a 4.00×10−2 M HF solution.
6) Find the percent dissociation of a 4.00×10−2 M HF solution.
1
Expert's answer
2020-04-06T12:06:23-0400

Find the pH and percent ionization of each HF solution. (Ka for HF is 6.8×10^−4.)

1) Find the pH of a 0.250 M HF solution

"[H^+] = \sqrt (C * Ka)"

"[H^+] = \sqrt (0.250 * 6.8*10^{-4}) = 1.3*10^{-2}"

"pH = -log(H^+)"

"pH = -log(1.3*10^{-2}) = 1.9"


2) Find the percent dissociation of a 0.250 M HF solution.

"\Alpha = ([H^+]/[HF])*100\%"

"\Alpha = (1.3*10^{-2}/0.250)*100\% = 5.2\%"


3) Find the pH of a 0.130 M HF solution.

"[H^+] = \sqrt (0.130 * 6.8*10^{-4}) = 9.4*10^{-3}"

"pH = -log(9.4*10^{-3}) = 2.03"


4) Find the percent dissociation of a 0.130 M HF solution.

"\Alpha = (9.4*10^{-3}/0.130)*100\% = 7.2\%"


5) Find the pH of a 4.00×10−2 M HF solution.

"[H^+] = \sqrt (4.00*10^{-2} * 6.8*10^{-4}) = 5.2*10^{-3}"

"pH = -log(5.2*10^{-3}) = 2.28"


6) Find the percent dissociation of a 4.00×10−2 M HF solution.

"\Alpha = (5.2*10^{-3}/4.0*10^{-2})*100\% = 13\%"